Complete the table. y = f(g(x)) u = g(x) y = f(u) = (5x – 875 U = y = 05

Complete the table. y = f(g(x)) u = g(x) y = f(u) y = (In x) 3 U= ly=u3
Answer #1

y = f (90) (6 = n y=flu) y = (5x-8) u=150-8 y=us (0)6in y=f(u) y = f(g(n) E a)? y=(en you? u=lenn